# PECCARY: A novel approach for characterizing orbital complexity, stochasticity, and regularity

### v1

@@ Abstract
Permutation Entropy and statistiCal Complexity Analysis for astRophYsics (PECCARY) is a computationally inexpensive, statistical method by which any time-series can be characterized as predominately regular, complex, or stochastic. Elements of the PECCARY method have been used in a variety of physical, biological, economic, and mathematical scenarios, but have not yet gained traction in the astrophysical community. This study introduces the PECCARY technique with the specific aims to motivate its use in and optimize it for the analysis of astrophysical orbital systems. PECCARY works by decomposing a time-dependent measure, such as the x-coordinate or orbital angular momentum time-series, into ordinal patterns. Due to its unique approach and statistical nature, PECCARY is well-suited for detecting preferred and forbidden patterns (a signature of chaos), even when the chaotic behavior is short-lived or when working with a relatively short duration time-seriesor small sets of time-series data. A variety of examples are used to demonstrate the capabilities of PECCARY. These include mathematical examples (sine waves, varieties of noise, sums of sine waves, well-known chaotic functions), a double pendulum system, and astrophysical tracer particle simulations with potentials of varying intricacies. Since the adopted timescale used to diagnose a given time-series can affect the outcome, a method is presented to identify an ideal sampling scheme, constrained by the overall duration and the natural timescale of the system. The accompanying PECCARY Python package and its usage are discussed.
Theoretical techniques(2093) — Galaxy dynamics(591) – Orbits(1184) — Orbit determination(1175) — Time series analysis(1916) — Exoplanet dynamics(490)
journal: ApJsoftware: galpy (Bovy, 2015)
@@ 1 Introduction
Permutation Entropy and statistiCal Complexity Analysis for astRophYsics (PECCARY) is a statistical method used to characterize a time-series as regular, stochastic (i.e., random or noisy), or complex, and identify its relevant timescales (Bandt & Pompe, 2002; Rosso et al., 2007; Weck et al., 2015). The use of Permutation Entropy and Statistical Complexity measures has been gaining traction in a wide variety of physical, biological, and mathematical scenarios, including plasma turbulence (Maggs & Morales, 2013; Gekelman et al., 2014; Weck et al., 2015; Maggs et al., 2015; Zhu et al., 2017), Solar wind and space plasma (Suyal et al., 2012; Weck et al., 2015; Ribeiro et al., 2017; Olivier et al., 2019; Weygand & Kivelson, 2019; Good et al., 2020), geological processes (Donner et al., 2015), river flow (Serinaldi et al., 2014; Thaxton et al., 2018), economic trends (Zunino et al., 2010; Bariviera et al., 2013; Araujo et al., 2020), biological or medical rhythms (Jordan et al., 2008; Li et al., 2010; Aronis et al., 2018), and for understanding the spread of the COVID-19 virus (Fernandes et al., 2020).
@@
One of the drivers of the evolution of a dynamical orbital system depends on the relative fraction and distribution of regular and complex orbits, as well as the timescales associated with each. Indeed, the formulation of chaos theory itself is firmly rooted in the study of chaotic behavior in astrophysical dynamical systems (e.g., the three body problem, Poincaré, 1891) and continues to inform studies of the secular evolution of galaxies (e.g., Fux, 2001; Pichardo et al., 2003; Patsis, 2006; Manos & Athanassoula, 2011; Valluri et al., 2016), planetary systems (e.g., Malhotra, 1993; Gladman, 1993; Saha & Tremaine, 1993; Astakhov et al., 2003; Lithwick & Wu, 2011; Deck et al., 2013), and black hole dynamics (e.g., Contopoulos, 1990; Suzuki & Maeda, 2000; Merritt & Poon, 2004), to name a few. The role of chaos in the evolution of dynamical systems is not the same as that of stochastic processes (whether physical or computationally induced, see Pfenniger, 1986; Kandrup & Willmes, 1994; Murray-Clay & Chiang, 2006; Sellwood & Debattista, 2009; Neyrinck et al., 2022, for various treatments), though they can be nearly indistinguishable in practice (Rosso et al., 2007) and quite often both are confusingly labeled “stochastic.” Differentiating between the two is particularly relevant in understanding the difference between dynamical processes and issues that arise from the limited resolution in discretized computation, like shot noise (e.g., Schaap & van de Weygaert, 2000; Dehnen, 2001; Varadi et al., 2003; Sellwood & Debattista, 2009; Sellwood, 2014).
@@
The PECCARY method is able to discern the nature of fluctuations in a time-series through its decomposition into a distribution of the occurrence frequency of patterns (described in Section 2.1). PECCARY can be set apart from well-known methods for determining regions of orbital chaos or irregularity, such as Lyapanov exponential divergence (e.g., Pichardo et al., 2003), Kolmogorov-Arnold-Moser (KAM) theory analysis (Weinberg, 2015a, b), Frequency Map Analysis (e.g., Laskar et al., 1992; Papaphilippou & Laskar, 1996, 1998; Valluri & Merritt, 1998; Valluri et al., 2012; Price-Whelan et al., 2016; Beraldo e Silva et al., 2019), and Surface of Section (SoS) analysis (e.g. Martinet, 1974; Athanassoula et al., 1983) since it is optimized to identify chaotic behavior on relatively short timescales and is agnostic to underlying physics. PECCARY expands the toolbox that astrophysical researchers have at their disposal for understanding dynamical systems. It provides an analysis technique that is suitable in situations where long-standing traditional techniques may not be as applicable, such as simulations with time-dependent potentials (e.g., a slowing bar or in systems that are accreting mass).
@@ 2 Overview of the PECCARY Method
PECCARY is comprised of two different statistical measures: Permutation Entropy and Statistical Complexity. The Permutation Entropy and Statistical Complexity measures were developed in the early 2000s (e.g., Bandt & Pompe, 2002; Rosso et al., 2007) as a way to distinguish noise from discrete chaotic maps, such as the logistic map or bifurcation diagram.
@@ 2.1 Determination of Ordinal Patterns
In their pioneering work, Bandt & Pompe (2002) developed the Permutation Entropy ($H$) measure as a means to identify chaotic behavior. Their approach relied on what they called “ordinal patterns.” An ordinal pattern is defined as the order in which a subset of $n$ sequential, discrete measurements from a given time-series appears such that their values increase from lowest relative amplitude to highest relative amplitude. In cases where there exist two equal values, the original order of points in the time-series is preserved. The values themselves are irrelevant since the magnitude of change between steps plays no role in this analysis.
@@
Ordinal patterns are extracted through a method that uses two parameters: the sampling size $n$ and the sampling interval $\ell$. The sampling size $n$ is the number of sequential points extracted to construct the ordinal patterns. Any time-series can be decomposed into consecutive, overlapping sets of $n$ time-steps, where the number of possible permutation orders is $n!$. The sampling interval $\ell$ is the integer number of timesteps from one sampled point to the next and probes the corresponding physical timescale in PECCARY (e.g., Zunino et al., 2012; Gekelman et al., 2014; Weck et al., 2015). The timescale for an extracted ordinal pattern associated with a given sampling size and sampling interval (known as the “pattern timescale”) is given by,
@@
Note that Bandt & Pompe (2002) called the sampling size the “embedding dimension” and the sampling interval the “embedding delay.” These refer to the same parameters, but this paper adopts different language for a more intuitive framing.
In order for PECCARY to produce meaningful results, the sampling size $n$ must be large enough that the set of possible ordinal patterns can be robustly used to describe the time-series ($n{>}2$, Bandt & Pompe, 2002), but not so large that the number of patterns becomes intractable. To the second point, PECCARY requires the condition $n!\ll N$ be met, where $N$ is the total number of sequential points in the time-series (Rosso et al., 2007). Bandt & Pompe (2002) noted that a practical choice lies between $3\leq n\leq 7$. They did not attempt a thorough proof; rather, they showed that chaotic behavior is well identified (even for noisy data) using values of $n$ within these bounds, with a slightly clearer signal near $n=6$. A sampling size of $n=5$ is adopted throughout this study, following the practice of several studies that effectively use Permutation Entropy (e.g., Cao et al., 2004; Rosso et al., 2007; Weck et al., 2015). A more thorough theoretical treatment is beyond the scope of this work but would yield useful justification for one’s choice of $n$ in future studies.
Figure 1: An arbitrary time-series of $N=19$ discrete points used to construct embedding dimension $n=5$ ordinal patterns. Shaded regions in panel (a) indicate three examples of patterns using sampling interval $\ell=1$. Shaded points in panels (b)-(d) show the ordinal pattern from lowest value (bottom) to highest value (top) of the five points within each shaded region. This ordinal pattern can be represented by the numerical sequence listed at the bottom of each panel.
@@ 2.2 Pattern Probability and Pattern Probability Distributions
After extracting the $\ell(n-1)$ ordinal patterns from the time-series, the pattern probability distribution $P$ (also called the “pattern distribution” by Weck et al. (2015)) can be produced for the $n!$ possible patterns. The probability $p(\pi_{i})$ for each pattern $\pi_{i}$ in $P$ is found by normalizing the occurrence frequency of that pattern so that
@@ 2.3.1 Permutation Entropy
A common metric for a pattern probability distribution $P$ is the Shannon Entropy (or information entropy) (Shannon, 1948), expressed as,
@@
is the Permutation Entropy (Bandt & Pompe, 2002; Rosso et al., 2007). Using this metric, a time-series that is dominated by a single pattern (e.g., a linear ramp), would have $H=0$, while an equally probable distribution of patterns (e.g., white noise, as in Figure 3(a)), would have $H=1$. An intermediate pattern, as in Figure 3(b), would have an intermediate value for $H$.
@@ 2.3.2 Disequilibrium and Complexity
where $S[P+P_{e}]$ is the Shannon Entropy for the sum of pattern probability distributions $P$ and $P_{e}$. The value of the disequilibrium $d(P,P_{e})$ is normalized by its maximum possible value ($d/d_{\rm max}$) is given by (Lamberti et al., 2004),
@@
and scales in the opposite direction to the Permutation Entropy (e.g., $D\rightarrow 1$ as $H\rightarrow 0$). Statistical Complexity, also known as the Jensen-Shannon Complexity, is given by the product (Lamberti et al., 2004; Rosso et al., 2007),
@@
The Statistical Complexity ($C$) measure has been compared to several established (e.g., Lyapanov analysis) and emerging (e.g., the LMC measure López-Ruiz et al., 1995) methods for identifying chaotic or complex behavior and is shown to be robust for a wide range of scenarios (Lamberti et al., 2004; Rosso et al., 2007), including logistic maps, the skew tent map, the Hénon map, the Lorenz map of Rossler’s oscillator, and Schuster maps (Schuster, 1988). Further, $C$ is an intensive measure that can be used to provide insight into the dynamics of a system (Lamberti et al., 2004), such as relevant timescales (explored in Sections 2.4 & 4). It is also reliably able to quantify the degree of chaos in systems that also have some degree of periodicity (Lamberti et al., 2004) or stochasticity (Rosso et al., 2007; Zunino et al., 2012).
@@ 2.4 The HC-Plane
Any time-series can be qualitatively sorted into its degree of regular, stochastic, and/or complex behavior by combining the metrics for the Permutation Entropy $H$ (normalized Shannon Entropy) and Statistical Complexity $C$ (normalized Jensen-Shannon Complexity) for a given sampling interval $\ell$ and sampling size $n$ (Rosso et al., 2007). Specifically, $H$ provides a quantitative scale for how stochastic or noisy the time-series is, while $C$ measures the degree of complexity or chaos by how many statistically preferred and/or forbidden patterns there are.
These regimes can be visualized as locations on a coordinate plane where the Permutation Entropy ($H$) is on the x-axis and the Statistical Complexity ($C$) is on the y-axis. Figure 4 illustrates the $HC$-plane with maximum and minimum limiting values for $C(H)$ indicated with solid curves, where the grey regions outside these curves are forbidden. The bounding curves are computed following a technique from  Calbet & Lopez-Ruiz (2001) using a Lagrange multiplier technique for each fixed Permutation Entropy from Equation 10. Regular time-series generate coordinates that occupy the left-hand region of the $HC$-plane while noisy time-series occupy the lower right. Complex or chaotic time-series occupy the upper middle region (Rosso et al., 2007; Zunino et al., 2012; Gekelman et al., 2014; Weck et al., 2015). Pure periodic functions (see discussion in §4.1.1) and circular orbits (see §4.3.1) fall on or within the region bounded by the dashed boundary line.
@@
There are three example time-series plotted on the $HC$-plane in Figure 4. These time-series are shown in Figure 3, where sampling parameters of sampling interval $\ell=1$ and sampling size $n=5$ were used to produce each $[H,C]$ coordinate in Fig. 4. The time-series generated from a uniform random number generator shown in Figure 3(a) (purple) is labeled as white noise and occupies the most extreme lower-right position in the $HC$-plane. The Hénon Map is a system described by (Hénon, 1976),

### v3

@@ Abstract
Permutation Entropy and statistiCal Complexity Analysis for astRophYsics (PECCARY) is a computationally inexpensive, statistical method by which any time-series can be characterized as predominantly regular, complex, or stochastic. Elements of the PECCARY method have been used in a variety of physical, biological, economic, and mathematical scenarios, but have not yet gained traction in the astrophysical community. This study introduces the PECCARY technique with the specific aims to motivate its use in and optimize it for the analysis of astrophysical orbital systems. PECCARY works by decomposing a time-dependent measure, such as the x-coordinate or orbital angular momentum time-series, into ordinal patterns. Due to its unique approach and statistical nature, PECCARY is well-suited for detecting preferred and forbidden patterns (a signature of chaos), even when the chaotic behavior is short-lived or when working with a relatively short duration time-series or small sets of time-series data. A variety of examples are used to demonstrate the capabilities of PECCARY. These include mathematical examples (sine waves, varieties of noise, sums of sine waves, well-known chaotic functions), a double pendulum system, and astrophysical tracer particle simulations with potentials of varying intricacies. Since the adopted timescale used to diagnose a given time-series can affect the outcome, a method is presented to identify an ideal sampling scheme, constrained by the overall duration and the natural timescale of the system. The accompanying PECCARY Python package and its usage are discussed.
Theoretical techniques (2093), Galaxy dynamics (591), Orbits (1184), Orbit determination (1175), Time series analysis (1916), Astronomical methods (1043), Astronomy software (1855)
journal: ApJsoftware: galpy (J. Bovy, 2015), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Numpy (C. R. Harris et al., 2020), Matplotlib (J. D. Hunter, 2007), SciPy (P. Virtanen et al., 2020)
show]soleyhyman@arizona.edu
@@ 1 Introduction
Permutation Entropy and statistiCal Complexity Analysis for astRophYsics (PECCARY) is a statistical method used to characterize a time-series as regular, stochastic (i.e., random or noisy), or complex, and identify its relevant timescales (C. Bandt & B. Pompe, 2002; O. A. Rosso et al., 2007; P. J. Weck et al., 2015). The use of Permutation Entropy and Statistical Complexity measures has been gaining traction in a wide variety of physical, biological, and mathematical scenarios, including plasma turbulence (J. E. Maggs & G. J. Morales, 2013; W. Gekelman et al., 2014; P. J. Weck et al., 2015; J. E. Maggs et al., 2015; Z. Zhu et al., 2017), Solar wind and space plasma (V. Suyal et al., 2012; P. J. Weck et al., 2015; H. V. Ribeiro et al., 2017; C. P. Olivier et al., 2019; J. M. Weygand & M. G. Kivelson, 2019; S. W. Good et al., 2020), geological processes (R. V. Donner et al., 2015), river flow (F. Serinaldi et al., 2014; C. S. Thaxton et al., 2018), economic trends (L. Zunino et al., 2010; A. F. Bariviera et al., 2013; F. H. A. d. Araujo et al., 2020), biological or medical rhythms (D. Jordan et al., 2008; D. Li et al., 2010; K. N. Aronis et al., 2018), and for understanding the spread of the COVID-19 virus (L. H. Fernandes et al., 2020).
@@
One of the drivers of the evolution of a dynamical orbital system depends on the relative fraction and distribution of regular and complex orbits, as well as the timescales associated with each. Indeed, the formulation of chaos theory itself is firmly rooted in the study of chaotic behavior in astrophysical dynamical systems (e.g., the three body problem, H. Poincaré, 1891) and continues to inform studies of the secular evolution of galaxies (e.g., R. Fux, 2001; B. Pichardo et al., 2003; P. A. Patsis, 2006; T. Manos & E. Athanassoula, 2011; M. Valluri et al., 2016), planetary systems (e.g., R. Malhotra, 1993; B. Gladman, 1993; P. Saha & S. Tremaine, 1993; S. A. Astakhov et al., 2003; Y. Lithwick & Y. Wu, 2011; K. M. Deck et al., 2013), and black hole dynamics (e.g., G. Contopoulos, 1990; S. Suzuki & K.-I. Maeda, 2000; D. Merritt & M. Y. Poon, 2004), to name a few. The role of chaos in the evolution of dynamical systems is not the same as that of stochastic processes (whether physical or computationally induced, see D. Pfenniger, 1986; H. E. Kandrup & D. E. Willmes, 1994; R. A. Murray-Clay & E. I. Chiang, 2006; J. A. Sellwood & V. P. Debattista, 2009; M. Neyrinck et al., 2022, for various treatments), though they can be nearly indistinguishable in practice (O. A. Rosso et al., 2007) and quite often both are confusingly labeled “stochastic.” Differentiating between the two is particularly relevant in understanding the difference between dynamical processes and issues that arise from the limited resolution in discretized computation, like shot noise (e.g., W. E. Schaap & R. van de Weygaert, 2000; W. Dehnen, 2001; F. Varadi et al., 2003; J. A. Sellwood & V. P. Debattista, 2009; J. A. Sellwood, 2014).
@@
The PECCARY method is able to discern the nature of fluctuations in a time-series through its decomposition into a distribution of the occurrence frequency of patterns (described in Section 2.1). PECCARY can be set apart from well-known methods for determining regions of orbital chaos or irregularity, such as Lyapanov exponential divergence (e.g., B. Pichardo et al., 2003), Kolmogorov-Arnold-Moser (KAM) theory analysis (M. D. Weinberg, 2015a, b), Frequency Map Analysis (e.g., J. Laskar et al., 1992; Y. Papaphilippou & J. Laskar, 1996, 1998; M. Valluri & D. Merritt, 1998; M. Valluri et al., 2012; A. M. Price-Whelan et al., 2016; L. Beraldo e Silva et al., 2019), and Surface of Section (SoS) analysis (e.g. L. Martinet, 1974; E. Athanassoula et al., 1983) since it is optimized to identify chaotic behavior on relatively short timescales and is agnostic to underlying physics. PECCARY expands the toolbox that astrophysical researchers have at their disposal for understanding dynamical systems. It provides an analysis technique that is suitable in situations where long-standing traditional techniques may not be as applicable, such as simulations with time-dependent potentials (e.g., a slowing bar or in systems that are accreting mass).
@@ 2 Overview of the PECCARY Method
PECCARY is comprised of two different statistical measures: Permutation Entropy and Statistical Complexity. The Permutation Entropy and Statistical Complexity measures were developed in the early 2000s (e.g., C. Bandt & B. Pompe, 2002; O. A. Rosso et al., 2007) as a way to distinguish noise from discrete chaotic maps, such as the logistic map or bifurcation diagram.
@@ 2.1 Determination of Ordinal Patterns
In their pioneering work, C. Bandt & B. Pompe (2002) developed the Permutation Entropy ($H$) measure as a means to identify chaotic behavior. Their approach relied on what they called “ordinal patterns.” An ordinal pattern is defined as the order in which a subset of $n$ sequential, discrete measurements from a given time-series appears such that their values increase from lowest relative amplitude to highest relative amplitude. In cases where there exist two equal values, the original order of points in the time-series is preserved. The values themselves are irrelevant since the magnitude of change between steps plays no role in this analysis.
@@
Ordinal patterns are extracted through a method that uses two parameters: the sampling size $n$ and the sampling interval $\ell$. The sampling size $n$ is the number of sequential points extracted to construct the ordinal patterns. Any time-series can be decomposed into consecutive, overlapping sets of $n$ time-steps, where the number of possible permutation orders is $n!$. The sampling interval $\ell$ is the integer number of timesteps from one sampled point to the next and probes the corresponding physical timescale in PECCARY (e.g., L. Zunino et al., 2012; W. Gekelman et al., 2014; P. J. Weck et al., 2015). The timescale for an extracted ordinal pattern associated with a given sampling size and sampling interval (known as the “pattern timescale”) is given by,
@@
Note that C. Bandt & B. Pompe (2002) called the sampling size the “embedding dimension” and the sampling interval the “embedding delay.” These refer to the same parameters, but this paper adopts different language for a more intuitive framing.
In order for PECCARY to produce meaningful results, the sampling size $n$ must be large enough that the set of possible ordinal patterns can be robustly used to describe the time-series ($n{>}2$, C. Bandt & B. Pompe, 2002), but not so large that the number of patterns becomes intractable. To the second point, PECCARY requires the condition $n!\ll N$ be met, where $N$ is the total number of sequential points in the time-series (O. A. Rosso et al., 2007). C. Bandt & B. Pompe (2002) noted that a practical choice lies between $3\leq n\leq 7$. They did not attempt a thorough proof; rather, they showed that chaotic behavior is well identified (even for noisy data) using values of $n$ within these bounds, with a slightly clearer signal near $n=6$. A sampling size of $n=5$ is adopted throughout this study, following the practice of several studies that effectively use Permutation Entropy (e.g., Y. Cao et al., 2004; O. A. Rosso et al., 2007; P. J. Weck et al., 2015). A more thorough theoretical treatment is beyond the scope of this work but would yield useful justification for one’s choice of $n$ in future studies.
Figure 1: An arbitrary time-series of $N=19$ discrete points used to extract ordinal patterns using sampling size $n=5$. Shaded regions in panel (a) indicate three examples of patterns using sampling interval $\ell=1$. Shaded points in panels (b)-(d) show the extracted points with their index numbers labeled on their markers. The corresponding ordinal pattern is listed at the top of each panel.
@@ 2.2 Pattern Probability and Pattern Probability Distributions
After extracting the $\ell(n-1)$ ordinal patterns from the time-series, the pattern probability distribution $P$ (also called the “pattern distribution” by P. J. Weck et al. (2015)) can be produced for the $n!$ possible patterns. The probability $p(\pi_{i})$ for each pattern $\pi_{i}$ in $P$ is found by normalizing the occurrence frequency of that pattern so that
@@ 2.3.1 Permutation Entropy
A common metric for a pattern probability distribution $P$ is the Shannon Entropy (or information entropy) (C. E. Shannon, 1948), expressed as,
@@
is the Permutation Entropy (C. Bandt & B. Pompe, 2002; O. A. Rosso et al., 2007). Using this metric, a time-series that is dominated by a single pattern (e.g., a linear ramp), would have $H=0$, while an equally probable distribution of patterns (e.g., white noise, as in Figure 3(a)), would have $H=1$. An intermediate pattern, as in Figure 3(b), would have an intermediate value for $H$.
@@ 2.3.2 Disequilibrium and Complexity
where $S[P+P_{e}]$ is the Shannon Entropy for the sum of pattern probability distributions $P$ and $P_{e}$. The value of the disequilibrium $d(P,P_{e})$ is normalized by its maximum possible value ($d/d_{\rm max}$) is given by (P. W. Lamberti et al., 2004),
@@
and scales in the opposite direction to the Permutation Entropy (e.g., $D\rightarrow 1$ as $H\rightarrow 0$). Statistical Complexity, also known as the Jensen-Shannon Complexity, is given by the product (P. W. Lamberti et al., 2004; O. A. Rosso et al., 2007),
@@
The Statistical Complexity ($C$) measure has been compared to several established (e.g., Lyapanov analysis) and emerging (e.g., the LMC measure R. López-Ruiz et al., 1995) methods for identifying chaotic or complex behavior and is shown to be robust for a wide range of scenarios (P. W. Lamberti et al., 2004; O. A. Rosso et al., 2007), including logistic maps, the skew tent map, the Hénon map, the Lorenz map of Rossler’s oscillator, and Schuster maps (H. G. Schuster, 1988). Further, $C$ is an intensive measure that can be used to provide insight into the dynamics of a system (P. W. Lamberti et al., 2004), such as relevant timescales (explored in Sections 2.4 & 4). It is also reliably able to quantify the degree of chaos in systems that also have some degree of periodicity (P. W. Lamberti et al., 2004) or stochasticity (O. A. Rosso et al., 2007; L. Zunino et al., 2012).
@@ 2.4 The HC-Plane
Any time-series can be qualitatively sorted into its degree of regular, stochastic, and/or complex behavior by combining the metrics for the Permutation Entropy $H$ (normalized Shannon Entropy) and Statistical Complexity $C$ (normalized Jensen-Shannon Complexity) for a given sampling interval $\ell$ and sampling size $n$ (O. A. Rosso et al., 2007). Specifically, $H$ provides a quantitative scale for how stochastic or noisy the time-series is, while $C$ measures the degree of complexity or chaos by how many statistically preferred and/or forbidden patterns there are.
These regimes can be visualized as locations on a coordinate plane where the Permutation Entropy ($H$) is on the x-axis and the Statistical Complexity ($C$) is on the y-axis. Figure 4 illustrates the $HC$-plane with maximum and minimum limiting values for $C(H)$ indicated with solid curves, where the grey regions outside these curves are forbidden. The bounding curves are computed following a technique from  Calbet & R. Lopez-Ruiz (2001) using a Lagrange multiplier technique for each fixed Permutation Entropy from Equation 10. Regular time-series generate coordinates that occupy the left-hand region of the $HC$-plane while noisy time-series occupy the lower right. Complex or chaotic time-series occupy the upper middle region (O. A. Rosso et al., 2007; L. Zunino et al., 2012; W. Gekelman et al., 2014; P. J. Weck et al., 2015). Pure periodic functions (see discussion in §4.1.1) and circular orbits (see §4.3.1) fall on or within the region bounded by the dashed boundary line.
@@
There are three example time-series plotted on the $HC$-plane in Figure 4. These time-series are shown in Figure 3, where sampling parameters of sampling interval $\ell=1$ and sampling size $n=5$ were used to produce each $[H,C]$ coordinate in Fig. 4. The time-series generated from a uniform random number generator shown in Figure 3(a) (purple) is labeled as white noise and occupies the most extreme lower-right position in the $HC$-plane. The Hénon Map is a system described by (M. Hénon, 1976),